OpenStudy (anonymous):

The ratio of the areas of two rectangles is 3 to 5. If the area of the larger rectangle is 85 square feet, what is the area of the smaller rectangle?

OpenStudy (anonymous):

51

OpenStudy (anonymous):

How do you set up and solve the proportion??

OpenStudy (anonymous):

3/5=x/85 x=85/5(3)=17... hve u got now

OpenStudy (anonymous):

I got it! Thank you! (=

OpenStudy (anonymous):

np..